OpenStudy (anonymous):

Solve equation for x, √4x+1-6=4

OpenStudy (anonymous):

where does the root end?

OpenStudy (anonymous):

On the one

sammixboo (sammixboo):

So is it \(\sqrt{4x+1-6}=4\)?

OpenStudy (anonymous):

Yes!

sammixboo (sammixboo):

OH, so So \(\sqrt{4x+1}-6=4\)?

sammixboo (sammixboo):

First or second one?

OpenStudy (anonymous):

Second :)

sammixboo (sammixboo):

OK. What do you think we do first?

OpenStudy (anonymous):

The 4x+1?

sammixboo (sammixboo):

Nope, the - 6 We do the opposite by adding 6 on both sides of the equation \(\sqrt{4x+1}-6+6 =4+6\) \(\sqrt{4x+1} =10\)

sammixboo (sammixboo):

Now we have to square both sides of the equation, remember? =)

OpenStudy (anonymous):

Okay

sammixboo (sammixboo):

\(4x+1=10^2\)

sammixboo (sammixboo):

Hold on a sec @sleepyjess can help! Got to go and do something real quick

OpenStudy (anonymous):

Okay, thank you!

sammixboo (sammixboo):

She's a really helpful user =3 No problem!

OpenStudy (sleepyjess):

awww thanks sammi :) Ok, so you 2 left off at 4x + 1 = \(10^2\)

OpenStudy (sleepyjess):

So what is 10^2?

OpenStudy (anonymous):

100?

OpenStudy (sleepyjess):

Yes! So after that we get 4x + 1 = 100. How do we get 1 to the right side?

OpenStudy (anonymous):

Would my answer be 99/4?

OpenStudy (sleepyjess):

Yes!

OpenStudy (anonymous):

Perfect! I understand it. Thank you both for the help!

OpenStudy (sleepyjess):

No problem! Thank you for participating :)